### v2

# MemeIntel: Explainable Detection of Propagandistic and Hateful Memes

Abstract
The proliferation of multimodal content on social media presents significant challenges in understanding and moderating complex, context-dependent issues such as misinformation, hate speech, and propaganda. While efforts have been made to develop resources and propose new methods for automatic detection, limited attention has been given to jointly modeling label detection and the generation of explanation-based rationales, which often leads to degraded classification performance when trained simultaneously. To address this challenge, we introduce MemeXplain, an explanation-enhanced dataset for propagandistic memes in Arabic and hateful memes in English, making it the first large-scale resource for these tasks. To solve these tasks, we propose a multi-stage optimization approach and train Vision-Language Models (VLMs). Our results show that this strategy significantly improves both label detection and explanation generation quality over the base model, outperforming the current state-of-the-art with an absolute improvement of $\sim{1.4}$% (Acc) on ArMeme and $\sim{2.2}$% (Acc) on Hateful Memes. For reproducibility and future research, we aim to make the MemeXplain dataset and scripts publicly available. [^1] https://github.com/MohamedBayan/MemeIntel
MemeIntel: Explainable Detection of Propagandistic and Hateful Memes
[…]
1 Introduction
Despite the rapid growth of multimodal content—integrating images, text, and sometimes video—the automated detection of harmful and false information on online news and social media platforms has become increasingly critical. In particular, identifying propaganda and hate in memes is essential for combating misinformation and minimizing online harm. While most research has focused on textual analysis, multimodal approaches have received comparatively less attention. In propaganda detection, text-based methods have evolved from monolingual to multilingual setups Piskorski et al. (2023); Hasanain et al. (2023), from binary to fine-grained span-level tasks Barrón-Cedeno et al. (2019); Habernal et al. (2017, 2018); Da San Martino et al. (2019). Hate speech detection has similarly progressed from text-based to multimodal approaches that integrate both textual and visual elements Kiela et al. (2020); Alam et al. (2022); Sharma et al. (2022).
Figure 1: Experimental steps for explanation generation and training.
The emergence of LLMs has demonstrated significant capabilities across various disciplines. To capture implicit hate and propaganda, LLMs utilize different techniques Cao et al. (2022); Kumar and Nandakumar (2022), such as prompt-based methods, contrastive learning, and cross-modal attention. Consequently, efforts have been made to leverage VLMs Zhang et al. (2024) to classify the harmful and propagandistic memes Cao et al. (2023). Despite progress, detecting implicit hate, especially with sarcasm or irony, remains challenging. Propagandistic memes add complexity through emotional appeals, humor, cultural cues, and manipulative language. These challenges are especially pronounced for English-centric multilingual models, which often struggle to capture nuances in non-English content due to cultural differences and linguistic variation. Specific challenges for Arabic often include cultural and political references that do not translate well into Western contexts, which makes it difficult for general-purpose models to detect accurately. For instance, memes about local conflicts or those using religious symbols often rely on subtle cues that are deeply rooted in cultural context. Arabic-focused datasets have made it clear that understanding these nuances is essential for accurate detection and meaningful explanations Alam et al. (2024b). To address these nuances, it is crucial for a system to provide not only accurate predictions but also clear, relatable explanations that align with the meme’s visual context and aid user understanding Hee et al. (2023); Yang et al. (2023); Huang et al. (2023); Sun et al. (2023).
An explanation-based approach offers numerous advantages and improves performance in various tasks Li et al. (2022); Magister et al. (2022); Nandi et al. (2024); Kumari et al. (2024). Although most studies have focused on textual content Li et al. (2022); Magister et al. (2022), some recent approaches Nandi et al. (2024); Kumari et al. (2024) have applied explainability to images. However, these methods rely on QA-based explanations that lack naturalness, use multiple inference with custom models, thus increasing computational complexity, and employ explanations only during training rather than as an inference output. To address these limitations, we propose a simplified method that delivers strong performance in meme classification and explanation generation across two datasets.
We summarize our main contributions below:
  • We developed an explanation-enhanced datasets, MemeXplain, using a rapid and low-cost annotation procedure;
  • We investigated state-of-the-art VLMs to identify an appropriate model for meme classification and explanation generation;
  • We proposed an efficient multi-stage optimization procedure that mitigates gradient conflicts and avoids catastrophic forgetting by formulating the optimization problem from the perspective of domain adaptation and task‑incremental learning.
  • We achieved state-of-the-art performance on two types of datasets related to propaganda and hateful content detection.
Our findings are as follows: (a) A higher human evaluation score suggests that explanations from stronger models (e.g., GPT-4o) are reliable and can serve as gold-standard explanations for training smaller models; (b) Task-specific fine-tuning improves performance over the base model; and (c) Our multi-stage optimization improves label detection and explanation quality by reducing gradient conflicts and avoiding catastrophic forgetting, achieving state-of-the-art performance. Overall, our work is the first to enhance VLMs for simultaneous propaganda and hateful content detection while providing natural reasoning to end users.
2 Related Work
The widespread use of social networks has become a major channel for spreading misinformation, propaganda, and harmful content. Significant research efforts have been directed toward addressing these challenges, particularly in multimodal disinformation detection Alam et al. (2022), harmful memes Sharma et al. (2022), and propagandistic content Dimitrov et al. (2021a). However, most studies have focused on detection, while less attention has been given to generating natural explanations/reasons behind the predicted labels.
Multimodal Propagandistic Content.
Following the previous research for propaganda detection using textual content Da San Martino et al. (2019), Dimitrov et al. (2021b) introduced SemEval-2021 Task 6 focusing on persuasion techniques detection in both textual and visual memes. Subsequently, the focus has extended to the detection of multilingual and multimodal propagandistic memes Dimitrov et al. (2024). Similar multimodal work on Arabic involves the development of datasets and shared task for propaganda detection Alam et al. (2024b); Hasanain et al. (2024). For the detection problem, typical approaches include a fusion of textual and visual embedding and a classification head on top them Hasanain et al. (2024); Shah et al. (2024), graph attention network based approach for multimodal objects Chen et al. (2024).
Multimodal Hate speech.
Similarly, there has been growing interest in detecting multimodal hate speech Kiela et al. (2020); Velioglu and Rose (2020); Hee et al. (2022). Due to the lack of resources, Kiela et al. (2020) developed a large-scale dataset for multimodal hate identification. This study advanced research in this area and emphasized the importance of integrating textual and visual features for effective detection. To further progress in this field, efforts have been made to develop resources for multiple languages, including Arabic Alam et al. (2024a), Bangla Hossain et al. (2022), and English Hee et al. (2023). A more detailed summary can be found in Sharma et al. (2022), which also highlights key challenges and outlines future research directions.
Training with Explanations.
Integrating reasoning or explainable capabilities to enhance LLM/VLM performance has been shown to be highly beneficial for various tasks across multiple domains Plaat et al. (2024). This approach has also proven effective for knowledge distillation and model compression Li et al. (2022); Magister et al. (2022), where explanations generated by large LLMs improve the performance and capabilities of smaller LLMs. For example, in the hate-speech detection task, Hare Yang et al. (2023) employs Chain-of-Thought (CoT) reasoning, while Huang et al. (2023) utilizes Chain of Explanation (CoE). Their aim is to improve the effectiveness of LLM-based sentiment classifiers by leveraging reasoning capabilities. Sun et al. (2023) proposed Clue and Reasoning Prompting (CARP), which uses keywords and reasoning to guide explanation.
CoT-Based Approaches.
CoT is a widely recognized prompting technique that generates a chain of reasoning to derive answers. Kumari et al. (2024) proposed a CoT-based framework for meme analysis using scene graphs to capture text- and image-based entity-object relationships. Their three-step prompting strategy guides the LLM to identify Emotion, Target, and Context for effective meme interpretation. SAFE-MEME Nandi et al. (2024) introduced two multimodal datasets and a CoT-based reasoning framework for meme-based hate speech detection, using Q&A-style prompts and hierarchical labels. However, their method was not evaluated on the Hateful Memes dataset, limiting direct comparison.
Comparison with Prior Work.
A key limitation of CoT-based methods is their reliance on multi-step reasoning, requiring multiple VLM inferences. In contrast, our approach: (a) avoids complex CoT steps, improving efficiency and reducing cost; and (b) generates explanations alongside classifications to enhance transparency and reliability. Hee et al. (2023) constructed a dataset providing explanations for hateful memes. However, unlike us, they focused solely on evaluating explanation generation and did not perform classification tasks. Despite the availability of their data, we do not use it due to the lack of naturalness. In particular, their explanations do not fully account for image content or image-centric contextual perspectives.
3 Dataset
3.1 ArMeme
The ArMeme dataset comprises approximately $\sim$6k manually annotated Arabic memes collected from various social media platforms Alam et al. (2024b). This dataset has been manually annotated with four labels such as Not propaganda, Propaganda, Not-meme and Other. Table 1 provides the distribution of the data splits. The memes with “Not propaganda” category covers over half of the dataset ($\sim$66%), followed by “Propaganda” and the distribution of “Not-meme“ and “Other“ classes are significantly smaller. This distribution highlights a substantial class imbalance, particularly between “Not propaganda” and the other categories.
Class label Train Dev Test Total
Not propaganda 2,634 384 746 3,764
Propaganda 972 141 275 1,388
Not-meme 199 30 57 286
Other 202 29 56 287
Total 4,007 584 1,134 5,725
Table 1: Data splits for ArMeme datasets.
3.2 Hateful Meme
The Hateful Memes dataset Kiela et al. (2020), is a benchmark designed to evaluate multimodal hate speech detection. It consists of $\sim$12k memes, combining both text and images, carefully curated to ensure that effective classification requires an understanding of both modalities. The dataset was created using a mix of synthetically generated memes and real-world examples, sourced from social media, while ensuring a balanced distribution of hateful and non-hateful content. In Table 2, we report the distribution of hateful meme dataset used for this study. Note that hateful meme dataset consists of two other splits (dev-seen and test-seen), here, we used unseen versions.
Class Label Train Dev-unseen Test-unseen Total
Not Hateful 5,481 340 1,250 7,071
Hateful 3,019 200 750 3,969
Total 8,500 540 2,000 11,040
Table 2: Distribution of hateful-meme dataset.
4 MemeXplain: Explanation Generation
The outcomes of an automatic system become more reliable for users if it provides decisions with adequate and interpretable natural explanations, which help users better understand the underlying reason behind the system’s decision Hee et al. (2023); Yang et al. (2023); Huang et al. (2023); Sun et al. (2023). Technically, this approach provides numerous advantages in terms of knowledge distillation, model compression, and enhancing the performance of target tasks in different domains Li et al. (2022); Magister et al. (2022); Nandi et al. (2024); Kumari et al. (2024). This motivates us to adopt the explanation-based approach in our research. However, we also aim to improve its efficiency, particularly with respect to dataset generation, model training, and system inference procedures.
In this research, we generate explanations for two different stages: (a) during existing dataset enhancement, which leverages an expert VLM (such as GPT-4o) to generate high-quality explanations and (b) during training/inference with a smaller VLM (such as Llama-3.2 11b). Mathematically, these two stages can be described by the functions $f(i,l)=e$ and $g(i)=(l,e)$, where $e$ denotes the explanation, $l$ is the label, and $i$ is the input image or meme (can also be associated text). Specifically, $f(i,l)$ is defined as the function that generates an explanation $e$ for dataset enhancement, given an image $i$ and its corresponding label $l$ as input; and $g(i)$ is defined as the function that generates a label-explanation pair $(l,e)$ during training and inference, given only the image content $i$ as input. In the data preparation phase, we use a VLM to generate explanations conditioned on both the input and the label using the function $f$. Next, in the training phase, the model learns to jointly predict the label and explanation from the input alone using the function $g$.
This research enhances two existing datasets with explanations, see Section 3 and Table 3 for the details and statistics. For the explanation generation task, it first uses a VLM for $f(i,l)$ and then involves human experts, which significantly accelerates high-quality explanation generation and lowers the overall cost and time. The following subsections provide step-by-step details.
Data
Total
Words
Avg.
Words
Total Expl.
Words
Avg.
Expl.
Words
Ar En Ar En
ArMeme
Train 58,688 15 280,341 375,843 70 94
Dev 8,583 15 40,756 55,336 70 95
Test 16,653 15 79,360 105,476 70 93
Total 83,924 15 400,457 536,655 70 94
Hateful Meme
Train 99,812 12 740,624 87
Dev 4,904 9 43,956 81
Test 18,079 9 173,982 87
Total 122,795 10 958,562 85
Table 3: Descriptive statistics of the dataset. Total Words and Avg. refer to the total and average number of words in the text. The last two columns represent the corresponding values for the explanations.
4.1 VLMs for Explanation Generation
Figure 1 illustrates an example of an Arabic meme along with its explanation-generation process using a VLM. We leverage GPT-4o (version 2024-11-20) for automated explanation generation. The choice of this model is motivated by prior studies Wang et al. (2023), which show that advanced GPT models can produce fluent, informative, persuasive, and logically sound explanations when properly prompted. In Listing 1, we present the prompts used for generating explanations for ArMeme and Hateful Memes. To refine the prompt, we iteratively tested several memes in both English and Arabic, selecting the one that produced the most reasonable explanations.
For Arabic memes, we generate two sets of explanations—one in English and one in Arabic. The motivation behind this approach is to assess the multilingual capability and quality of smaller VLMs, such as Llama-3.2 11b, in generating explanations and labels in both languages.
Size of the Explanation
Determining the optimal length for the explanations is important to balance informativeness and cognitive load Herm (2023). Shen et al. (2022) explored the relationship between the length of the explanation and human understanding, finding that the shortest rationales are often ineffective. Recently, Wang et al. (2023) also studied the effect of explanation size and found that human evaluators are less willing to read longer explanations. To achieve an optimal balance, we iteratively tested various explanation lengths and ultimately set a limit of 100 words.
Model and its Parameters
To utilize GPT-4o OpenAI (2023), we accessed the OpenAI API via Azure services. Although recently released o1 models have shown promising directions for complex reasoning, they were not accessible to us. For explanation generation, we used zero-shot inference. To ensure reproducibility, we set the temperature value to zero.
4.2 Manual Annotation of Explanation
Given that our idea is to use the generated explanation as gold data for further training and evaluation, therefore, we intended to go through human evaluation process. Following the prior studies Wang et al. (2023); Huang et al. (2023); Agarwal et al. (2024) we adopted four metrics discussed below. For each metric we use 5-point Likert scale.
Informativeness. Measures the extent to which the explanation provides relevant and meaningful information for understanding the reasoning behind the label. A highly informative explanation offers detailed insights that directly contribute to the justification, while a low-informative explanation may be vague, incomplete, or lacking key details.
Clarity. Assesses how clearly the explanation conveys its meaning. A clear explanation is well-structured, concise, and easy to understand without requiring additional effort. It should be free from ambiguity, overly complex language, or poor phrasing that might hinder comprehension.
Plausibility. Refers to the extent to which an explanation logically supports the assigned label and appears reasonable given the meme’s content. A plausible explanation should be coherent, factually consistent, and align with the expected reasoning behind the label.
Faithfulness. Measures how accurately an explanation reflects the reasoning behind the assigned label. A faithful explanation correctly represents the key factors and logical steps that justify the label, without adding misleading or unrelated details.
Annotation Setup. For manual annotation, we first prepared an annotation guideline and a platform (see Appendix B and A, respectively) for the annotators. For the Arabic memes, we recruited annotators who are native Arabic speakers and fluent in English, all holding at least a bachelor’s degree. Because of their fluency, they also handled the hateful meme. We provided necessary training and consultation, and all had prior experience with similar tasks. A total of six annotators participated in the evaluation. In line with institutional requirements, each signed a Non-Disclosure Agreement (NDA), and a third-party company managed their compensation at standard hourly rates based on their location.
Annotation Agreement for Explanation. In Table 4, we summarize the annotation agreement scores of the explanations. We used 5-point Likert scale for various human evaluation metrics, including informativeness, clarity, plausibility, and faithfulness. We compute the average Likert scale value (from three annotators) for all evaluation metrics. We manually evaluated complete test sets for both ArMeme and Hateful meme except ArMeme with English explanation. For the later case we randomly selected 200 random examples. This decision was made to reduce the human annotation cost. The average agreement scores for the ArMeme dataset with Arabic explanations are >4.5>4.5 out of 5 indicating high agreement in all evaluation metrics. However, for the English explanations of ArMeme, the faithfulness and plausibility scores are relatively lower. To better understand this issue, we plan to conduct further annotation on the complete test set of explanations. For the Hateful Memes dataset, the average Likert scale agreement scores range from 4.04 to 4.10 out of 5.
In addition, we also computed the annotation agreement on ordinal scales by adopting the agreement index $r^{*}_{wg(j)}$ James et al. (1984), which compares observed variance in ratings to the maximum possible variance under complete disagreement. For each item, the agreement score is computed as: $r^{*}_{wg(j)}=1-\frac{S_{X}^{2}}{\sigma^{2}_{\text{mv}}}$ where $S_{X}^{2}$ is the observed variance across annotators and $\sigma^{2}_{\text{mv}}$ is the maximum variance possible given the scale (computed as $\sigma^{2}_{\text{mv}}=0.5(X_{U}^{2}+X_{L}^{2})-[0.5(X_{U}+X_{L})]^{2}$, with $X_{U}=5$ and $X_{L}=1$ for a 5-point scale). The average agreement scores for ArMeme and Hateful memes are above 0.83 and 0.92, respectively, for all metrics. These values indicate strong agreement O’Neill (2017).
Dataset Faithfulness Clarity Plausiblity Informative
ArMeme Ar expl. 4.64 4.69 4.69 4.74
ArMeme En expl. 3.91 4.50 3.81 4.13
Hateful meme 4.01 4.18 4.04 4.10
Table 4: Average Likert scale value for each human evaluation (annotation) metric across different sets of explanations.
4.3 Basic Statistics
Table 3 presents the basic statistics for both datasets. The average explanation length is 94 words for Arabic and 85 words for English. Notably, we instructed GPT-4o to generate explanations with fewer than 100 words.
5 Methodology
5.1 Instructions Dataset
Our approach follows the standard pipeline for aligning LLMs with user intentions and specific tasks through fine-tuning on representative data Zhang et al. (2023); Kmainasi et al. (2024); Hasan et al. (2025). This process typically involves curating and constructing instruction datasets that guide the model’s behavior, ensuring it generates responses that align with the desired objectives. For our study, the responses include label and explanation. Hence, we created instruction format for both datasets. For the ArMeme dataset, we replicated the experiments for both Arabic and English explanations.
5.2 Model Selection
As shown in Figure 1, our first experimental phase involves model selection among several recent VLMs, including Llama-3.2 (11b) Dubey et al. (2024), Paligemma 2 (3b) Steiner et al. (2024), Qwen2-vl Wang et al. (2024), and Pixtral (12b) Agrawal et al. (2024). The goal here is to select the VLM that maximizes $p(l|i)$ for all image samples.
We evaluate the base models in a zero-shot setting and fine-tune them using an instruction-following paradigm. The instructions prompt the model to generate responses in the format ‘‘Label: (class_label)’’. We use a regex-based function to extract the predicted labels.
Note that this stage fine-tunes the models to predict class labels only, allowing us to verify whether they can handle multilingual inputs—especially in understanding Arabic text, cultural nuances, and image context. We do not ask the model to generate explanations here, as that is a more complex task and could affect their performance.
Based on the results reported in Tables 5 and 6, we selected Llama-3.2-vision-instruct (11b) for further training with explanations.
5.3 Multi-Stage (MS) Optimization Procedure
To emphasize our contribution, we introduce a novel optimization procedure to train VLM with MemeXplain, which decouples the classification and explanation generation tasks. While training both tasks in a single step may seem like an obvious choice, in practice, the two objectives produce conflicting gradient signals due to the fundamental differences between the learning objectives. Classification requires precise mapping of multimodal cues to discrete labels, whereas explanation generation demands fluency and coherence in free-form natural language. Merging them too early can compromise the model’s performance on both tasks. Conversely, training them completely separately risks overwriting knowledge of one task while learning the other. To address these challenges, we propose a two-stage optimization procedure that decouples the learning objectives by first optimizing in the classification domain in stage-1, followed by augmentation through explanation generation in stage-2. The overall goal of this optimization procedure is to obtain an optimal VLM that maximizes $p(l,e|i)$ for the training dataset (cf. Continual Training in Figure 1).
Problem Formulation.
We formulate the joint optimization problem as follows: $L_{\mathrm{total}}=L_{\mathrm{classif}}+W_{\mathrm{expl}}\;\cdot\;L_{\mathrm{expl}},$ where $L_{\mathrm{classif}}$ and $L_{\mathrm{expl}}$ are the classification and explanation losses, respectively. $W_{\mathrm{exp}}$ is a step‑function weight that switches from 0 to 1 between the stages. This formulation draws on the principles from Domain Adaptation and Task‑Incremental Learning (Van de Ven et al., 2022) with the aim to: (a) Isolate classification learning and avoid conflicting updates and (b) Prevent catastrophic forgetting by gradually integrating the explanation objective.
Stage 1 - Classification Fine-Tuning.
We set $W_{\mathrm{exp}}=0$ and optimize exclusively on $L_{\mathrm{classif}}$. It adapts the pretrained VLM to the domain of hateful/propagandistic content, establishing a strong feature backbone for accurate label prediction.
Stage 2 - Joint Classification & Explanation Fine‑Tuning.
We set $W_{\mathrm{exp}}=1$ to optimize $L_{\mathrm{classif}}+L_{\mathrm{explanation}}$. After obtaining the domain-adapted backbone from Stage 1, the model proceeds to learn how to generate coherent, contextually grounded explanations alongside accurate classifications. This stepwise integration ensures that the model preserves its classification capabilities and avoids catastrophic forgetting, while developing proficiency in natural language reasoning.
By decoupling and then recombining these objectives in a straightforward two-phase procedure, our multi-stage (MS) optimization presents an easily implementable extension to standard VLM training pipelines. To validate its effectiveness, we compare it against a single-stage (SS) fine-tuning baseline, where the model is directly trained on the label-with-explanation dataset. Our ablation studies (detailed in Section 6) demonstrate that the proposed multi-stage approach significantly outperforms the single-stage strategy.
Model Setup Acc (%) W-F1 M-F1
Alam et al. (2024b) Qarib 69.7 0.690 0.551
Alam et al. (2024b) mBERT 70.7 0.675 0.487
Llama-3.2 (11b) Base 13.4 0.172 0.113
Llama-3.2 (11b) FT 68.0 0.665 0.452
Paligemma2 (3b) Base 15.3 0.090 0.080
Paligemma2 (3b) FT 65.9 0.524 0.200
Qwen2V (7b) Base 63.1 0.550 0.242
Qwen2V (7b) FT 72.2 0.686 0.440
Pixtral (12b) Base 14.6 0.177 0.133
Pixtral (12b) FT 70.8 0.636 0.377
Table 5: Results for ArMeme. FT: Fine-tuned. Qarib Abdelali et al. (2021) is a Arabic BERT (text only). mBERT - multilingual BERT (text only).
Model Setup Acc (%) W-F1 M-F1
Kiela et al. (2020) 69.47±2.06
Cao et al. (2022) 72.98±1.09
Burbi et al. (2023) 77.70
Llama-3.2 (11b) Base 66.1 0.650 0.618
Llama-3.2 (11b) FT 77.7 0.770 0.748
Paligemma2 (3b) Base 35.2 0.277 0.217
Paligemma2 (3b) FT 69.2 0.664 0.623
Qwen2V (7b) Base 66.4 0.669 0.442
Qwen2V (7b) FT 77.9 0.773 0.753
Pixtral (12b) Base 66.7 0.667 0.430
Pixtral (12b) FT 77.2 0.766 0.746
Table 6: Results for Hateful meme. FT: Fine-tuned
5.4 Training Setup
Our fine-tuning experiments utilize QLoRA Dettmers et al. (2023), which combines INT4 quantization with parameter-efficient fine-tuning through Low-Rank Adaptation (LoRA) Hu et al. (2022). In our setup, the base model is quantized to 4-bit precision, with LoRA updates applied to a subset of the model parameters. This approach was selected to address computational and memory resource constraints. We adapted all relevant submodules (vision, language, attention, and MLP layers) with a LoRA rank of 16, an alpha of 16, and no dropout. We used a per-device batch size of 2 with gradient accumulation over 4 steps, optimizing with AdamW, a weight decay of 0.01, and a linear scheduler with 5 warmup steps.
Single-Stage (SS) Baseline.
We implement a baseline where models are fine-tuned end-to-end using a unified classification and classification-with-explanation objectives. Model selection is performed on the development set, and final results are reported on the test set using the best-performing checkpoint.
Multi-Stage (MS) Curriculum.
We design a staged curriculum to progressively instill multimodal reasoning capabilities into the model. In the first stage, the model is trained exclusively on the classification dataset, and intermediate checkpoints are evaluated on the development set. The best checkpoint from this stage is then used to initialize the second stage, where the model is further optimized with the classification-with-explanation objective. Hyperparameter exploration revealed that lowering the learning rate while increasing the number of training epochs in the second stage yields performance gains in our setup.
5.5 Evaluation Setup and Metrics
The results for pretrained models, fine-tuned models, and baselines are summarized in Tables 5 and 6. For performance measurement across different experimental settings, we compute accuracy, weighted F1, and macro-F1. We evaluate the model’s explanation performance on the test set using semantic similarity-based metric, measured by the F1 score within BERTScore (Zhang et al., 2020). This score is computed using contextual embeddings extracted from pre-trained BERT models. To enhance accuracy, we utilize language-specific transformer models for embedding extraction. For Arabic we use AraBERT (v2) Antoun et al. (2020) model and for English we use bert-base-uncased model Devlin et al. (2019). In addition, we also compute BLEU and METEOR scores.
6 Experimental Results and Discussion
This section first presents competitive results among our proposed method and the state-of-the-art approaches. Next, it briefly analyzes and investigates the proposed method to validate and highlight the core contributions of this research. To increase clarity, we emphasize that all the results discussed in this section are based on an experimental setup that evaluates models using the MemeXplain dataset, which includes both classification and explanation for training. This differs from the results presented in the previous section, in Tables 5 and 6, where we evaluated and compared models trained on the original datasets, which include only classification labels. Therefore, none of the results discussed in this section should be compared to those presented in Tables 5 and 6.
Comparison with State‑of‑the‑Art.
Table 7 presents the comparison. On the ArMeme dataset, our proposed approach (Llama MS) achieves the highest accuracy at 72.1% and the best weighted F1 score at 0.699, with Qarib and mBERT trailing behind. Although Qarib achieves the highest macro F1 score (0.551), Llama MS remains competitive with a macro F1 of 0.536. Importantly, our approach stands out by providing explanations that add significant value. On the Hateful Meme dataset, Llama MS surpasses the state-of-the-art by achieving the best performance: an accuracy of 79.9%, a weighted F1 score of 0.802, and a macro F1 score of 0.792. These results clearly highlight the advantages of our explainability-enhanced dataset and the effectiveness of our proposed optimization procedure for both classification and explanation-generation tasks.
Model Setup Acc(%) W-F1 M-F1
ArMeme
Alam et al. (2024b) Qarib 69.7 0.690 0.551
Alam et al. (2024b) mBERT 70.7 0.675 0.487
Alam et al. (2024b) ResNet50 66.0 0.637 0.434
Llama MS FT 72.1 0.699 0.536
Llama (Ar-Exp) MS FT 72.0 0.696 0.499
Hateful Meme
Kiela et al. (2020) 69.47±2.06
Cao et al. (2022) 72.98±1.09
Burbi et al. (2023) 77.70
Llama MS FT 79.9 0.802 0.792
Table 7: Comparison with SOTA and our results (Llama MS and Llama (Ar-Exp) MS). ResNet50 He et al. (2016) is an image only model. MS: Multi-stage (Our approach).
Model Setup Acc (%) W-F1 M-F1 BS BL M
ArMeme
Llama Base 12.7 0.165 0.105 0.61 0.24 0.17
Llama SS FT 68.2 0.584 0.257 0.70 0.56 0.36
Llama MS FT 72.1 0.699 0.536 0.70 0.57 0.35
Llama Ar-Exp Base 19.0 0.246 0.125 0.58 0.12 0.09
Llama MS Ar-Exp FT 72.0 0.696 0.499 0.72 0.55 0.29
Hateful Meme
Llama Base 65.2 0.615 0.567 0.661 0.35 0.23
Llama SS FT 75.9 0.760 0.745 0.767 0.65 0.47
Llama MS FT 79.9 0.802 0.792 0.777 0.67 0.49
Table 8: Results with ArMeme and Hateful meme classification and explanation generation. Llama: Llama-3.2 (11b), BS: BERTScore, BL: BLEU, M: METEOR. $S$: Single-stage, MS: Multi-stage. Ar-Exp: Model trained with Arabic explanation. W-F1: Weighted F1, M-F1: macro-F1
Ablation Study on Different Model Training Settings.
Table 8 presents classification and explanation-generation results on the ArMeme and Hateful Meme datasets across different settings: (a) Llama represents the performance of the base model without fine-tuning; (b) Llama SS represents the performance after single-stage fine-tuning; and (c) Llama MS represents the performance after completing the second stage of our proposed multi-stage optimization. The extension Ar-Exp after a setting name indicates that the model was trained using Arabic explanations. Based on these settings, we analyze the results from various perspectives: (a) Base vs. FT: demonstrates the performance difference between the same model with and without fine-tuning (FT);
(b) Single-stage (SS) vs. Multi-stage (MS): highlights the necessity and benefits of the proposed optimization procedure; and
(c) Eng-Exp vs. Ar-Exp: showcases the multilingual capability of the selected VLM.
First, we compare the Base vs. FT setup in Table 8, from which it is evident that the FT model significantly outperforms the baseline. For example, on the ArMeme dataset, while the baseline (Llama) achieves an accuracy of 12.7%, the proposed fine-tuning (Llama-MS) boosts it to 72.1%. Similarly, on the Hateful Meme dataset, fine-tuning improves the base accuracy from 65.2% (Llama) to 79.9% (Llama-MS). We observe similar improvements in the F1 metrics for classification and BERTScore for explanation quality. These significant performance gains validate our approach of fine-tuning the base models with the explainability enhanced dataset, demonstrating its efficacy for the meme classification and explanation generation tasks.
Next, we compare the SS vs. MS setup in Table 8, which reveals that multi-stage (MS) fine-tuning further enhances performance over the classification single-stage approach. For example, on the ArMeme dataset, the accuracy increased from 68.2% (Llama-SS) to 72.1% (Llama-MS), the weighted F1 increased from 0.584 (Llama-SS) to 0.699 (Llama-MS), the macro F1 increased significantly from 0.257 (Llama-SS) to 0.536 (Llama-MS), and the BERTScore for Arabic explanation increased significantly from 0.58 (Llama-SS) to 0.72 (Llama-MS). A similar trend is observed on the Hateful Meme dataset, where additional fine-tuning iterations yield more robust classification (approximately 4% improvement) and enhanced explanation quality. These performance gains validate our proposed multi-stage optimization procedure to further refine the VLMs.
Assessment of Multilingual Capability.
We compare Llama MS - FT with Llama MS Ar-Exp - FT in Table 8, which shows that fine-tuning using explanations generated in both languages yields comparable outcomes. This validates the multilingual capability of our empirically chosen VLM for the target task and helps users understand multilingual content without fluency in that language. For example, our model allows an English speaker to analyze Arabic memes and receive explanations in English.
Annotation Agreement: Additional Analysis
We report human evaluations of the generated explanations from two models: (i) Llama MS (ArMeme) and (ii) Llama MS (Hateful Memes), as detailed in Table 9. Due to resource constraints, the manual evaluation was conducted on a representative sample of 100 instances. For the ArMeme dataset, the scores across all evaluation metrics range from 4.15 to 4.74 out of 5, while for the Hateful Memes dataset, the scores range from 4.41 to 4.54, indicating consistently strong performance across both datasets.
Dataset Faithfulness Clarity Plausiblity Informative
ArMeme 4.63 4.74 4.56 4.15
Hateful meme 4.41 4.44 4.43 4.54
Table 9: Average Likert scale value for each human evaluation (annotation) metric for the explanations of different datasets.
Error Analysis.
To better understand the model’s reasoning capabilities and failure modes, we conduct an error analysis contrasting correct and incorrect label predictions, as well as comparing single-stage and multistage training paradigms. Our analysis reveals that while the model often aligns well with human explanations in correctly classified instances, it struggles with nuanced cases involving humor or implicit context. Furthermore, multistage training consistently outperforms single-stage training in label accuracy by better integrating textual and visual cues while providing explanation. See Section D for detailed examples and discussion.
7 Conclusions and Future Work
In this study, we introduce a MemeXplain dataset for propagandistic and hateful meme detection and natural explanation generation, making it the first resource of its kind. To address both detection and explanation generation tasks and ensure efficient VLMs model training on this dataset, we also propose a multi-stage optimization procedure. To evaluate the multilingual capability of the model, we developed Arabic and English explanations for Arabic memes. The inclusion of English explanations benefits non-Arabic speakers, whereas providing explanations in the native language ensures that cultural nuances are accurately conveyed. With our training procedure, we demonstrate improved detection performance for both ArMeme and hateful memes. The higher performance of explanation generation further demonstrates the efficacy of our multi-stage training approach.
Moreover, the proposed multi-stage scheme is agnostic to specific VLM architectures and applicable in contexts where tasks impose divergent gradient demands, thereby offering a versatile framework for multi-task vision–language learning.
We foresee several future directions to extend this research and explore the following: (a) training the model with additional data through data augmentation, which could help it become an instruction-generalized model and potentially enhance its performance further; (b) incorporating pseudo and self-labeled data using an active learning procedure to incrementally improve the model’s capabilities; and (c) developing a task-generalized model that addresses multiple tasks.
8 Limitations
Due to the complex nature of manual explanation creation, we have relied on GPT-4o for explanation generation. To ensure the reliability of the explanation we have manually evaluated in four criteria such as informativeness, clarity, plausiblity, and faithfulness on a small sample for each set of explanation. The preliminary evaluation scores suggest that we can rely on the gold explanation as the reference. As a part of ongoing work we plan to conduct manual evaluation on a larger set. An important aspect of the ArMeme dataset is that it is highly imbalanced, which affects overall performance. One possible approach to address this issue is to increase the number of memes labeled as propaganda, other, and not-meme. This can be achieved through data augmentation or by collecting additional memes.
Ethics and Broader Impact
We extended existing datasets by adding explanations. To the best of our knowledge, the dataset does not contain any personally identifiable information, making privacy risks nonexistent. Regarding the explanations, we provided clear annotation instructions and cautioned annotators that some memes might be offensive. It is important to note that annotations are inherently subjective, which can introduce biases into the overall evaluation results. We encourage researchers and users of this dataset to remain critical when developing models or conducting further research. Models built using this dataset could be highly valuable for fact-checkers, journalists, and social media platforms.
Acknowledgments
The work of F. Alam is supported by the NPRP grant 14C-0916-210015 from the Qatar National Research Fund part of Qatar Research Development and Innovation Council (QRDI). The findings achieved herein are solely the responsibility of the authors.
Appendix A Annotation Guideline
You will be shown a meme, a label assigned to it, and an explanation for the assigned label. As an annotator, your task is to carefully examine each meme, label, and explanation. Then assess the quality of the explanation provided for the assigned label. Follow the steps below to ensure a thorough evaluation:
Analyze the Meme
  • Observe the image and read the accompanying text.
  • Understand the overall message and the potential implications of the meme.
Check the Assigned Label
  • Check the given label. The label is the result of annotation done by multiple human annotators.
Evaluate the Explanation
  • Read the explanation provided for why the meme has been assigned its label.
  • Assess the explanation based on the metrics below. Each metric is scored on a Likert scale from 1-5.
Kindly note that to evaluate the explanation, you do not have to agree or disagree with the given label.
A.1 Metrics
A.1.1 Informativeness
Measures the extent to which the explanation provides relevant and meaningful information for understanding the reasoning behind the label. A highly informative explanation offers detailed insights that directly contribute to the justification, while a low-informative explanation may be vague, incomplete, or lacking key details.
As an annotator, you are judging if the explanation provides enough information to explain the label assigned to the meme.
  • 1 = Not informative: The explanation lacks relevant details and does not help understand why the meme is labeled as such.
  • 2 = Slightly informative: The explanation provides minimal information, but key details are missing or unclear.
  • 3 = Moderately informative: The explanation contains some useful details but lacks depth or supporting reasoning.
  • 4 = Informative: The explanation is well-detailed, providing a clear and meaningful justification for the label.
  • 5 = Very informative: The explanation is thorough, insightful, and fully justifies the label with strong supporting details.
A.1.2 Clarity
Assesses how clearly the explanation conveys its meaning. A clear explanation is well-structured, concise, and easy to understand without requiring additional effort. It should be free from ambiguity, overly complex language, or poor phrasing that might hinder comprehension.
As an annotator, you are judging the language and structure of the explanation. Spelling mistakes, awkward use of language, and incorrect translations will negatively impact this metric.
  • 1 = Very unclear: The explanation is confusing, vague, or difficult to understand.
  • 2 = Somewhat unclear: The explanation has some clarity but includes ambiguous or poorly structured statements.
  • 3 = Neutral: The explanation is somewhat clear but may require effort to fully grasp.
  • 4 = Clear: The explanation is well-structured and easy to understand with minimal ambiguity.
  • 5 = Very clear: The explanation is highly readable, precise, and effortlessly understandable.
A.1.3 Plausibility
Refers to the extent to which an explanation logically supports the assigned label and appears reasonable given the meme’s content. A plausible explanation should be coherent, factually consistent, and align with the expected reasoning behind the label. While it does not require absolute correctness, it should not contain obvious contradictions or illogical claims.
As an annotator, you are judging if the explanation actually supports the label assigned to the meme. For example, if a meme is labeled as Not Propaganda, the explanation given should justify that label.
  • 1 = Not plausible at all: The explanation does not align with the label and seems completely incorrect.
  • 2 = Weakly plausible: The explanation has some relevance but lacks strong justification or contains logical inconsistencies.
  • 3 = Moderately plausible: The explanation somewhat supports the label but may be incomplete or partially flawed.
  • 4 = Plausible: The explanation logically supports the label and is mostly reasonable.
  • 5 = Highly plausible: The explanation is fully aligned with the label and presents a strong, logical justification.
A.1.4 Faithfulness
Measures how accurately an explanation reflects the reasoning behind the assigned label. A faithful explanation correctly represents the key factors and logical steps that justify the label, without adding misleading or unrelated details. High faithfulness means the explanation stays true to the actual reasoning used for classification, ensuring reliability and consistency.
As an annotator, you are judging how well the explanation reflects the logic behind the label. For example, if the explanation claims an implication of the meme, it should also present the logical reasoning behind it.
  • 1 = Not faithful at all: The explanation is completely unrelated to the given label and does not reflect a valid reasoning process.
  • 2 = Weakly faithful: Some elements of the explanation are relevant, but much of it is misleading, inconsistent, or lacks proper justification.
  • 3 = Moderately faithful: The explanation captures parts of the reasoning but includes unrelated, unclear, or unnecessary justifications.
  • 4 = Faithful: The explanation aligns well with the reasoning behind the label and includes relevant, logical details.
  • 5 = Highly faithful: The explanation fully and accurately reflects the correct reasoning, without any misleading or irrelevant information.
Appendix B Annotation Platform
In Figure 2, we present the screenshot of the interface designed for the explanation evaluation of hateful meme, which consisted of an image, respective label, and explanation for the label, annotation guidelines, and four different evaluation metrics. We used 5-point Likert scale for each evaluation metric. Annotators select one of the Likert scale value following the annotation guideline for each metric and submit.
Figure 2: A screenshot of the annotation platform for the explanation evaluation of hateful meme.
Appendix C Prompt for Explanation Generation
In Listings 1 and 2, we provide the prompts used to generate explanations for ArMeme and Hateful Meme. The prompt in Listing 1 is specifically for generating Arabic explanations for ArMeme. To generate English explanations, the same prompt was used, except it was adapted for English.
You are a Propaganda Image Detection Expert. A human expert has already classified the image as {class_label}.
Do not change or re-identify the classified label of the image.
First, analyze the images visual elements (objects, symbols, color usage, composition)
and provide a concise description. Next, read and paraphrase any text in the image-especially if
it is in non-Arabic-so that a Arabic speaker could understand its literal or intended meaning.
Explain how that text, in conjunction with the visual elements, supports or illustrates the known
propaganda technique. Think carefully about emotional appeals, humor, cultural references, manipulative
language, or other rhetorical strategies.
Finally, provide a thoughtful explanation in Arabic (up to {explanation_length} words) of why these elements led the human to classify
the image as propaganda. Be specific about how the text and visuals interact, referencing the
Arabic context or cultural nuances if relevant. Your answer must be valid JSON with a single field:
"explanation": "Your explanation here (up to {explanation_length} words) in Arabic."
Listing 1: Prompt for generating Arabic explanation for ArMeme dataset. The place-holders refers to label, and number of words, respectively.
You are a Hateful Image Detection Expert. A human expert has already classified the image as {}. Do not change or re-identify the classified label of the image.
First, analyze the images visual elements (objects, symbols, color usage, composition) and provide a concise description. Next, read and paraphrase any text in the image so anyone speaking English can understand its literal or intended meaning.
Explain how that text, in conjunction with the visual elements, supports or illustrates the known hateful content.
Finally, provide a thoughtful explanation in English (up to {} words) of why these elements led the human to classify the image as hateful. Be specific about how the text and visuals interact, referencing the context or cultural nuances if relevant. Your answer must be valid JSON with a single field:
"explanation": "Your explanation here (up to {} words) in English."
Listing 2: Prompt for generating explanation. The place-holders refers to label, and number of words, respectively.
Appendix D Error Analysis
D.1 Correct vs. Incorrect Label Prediction
In Figure 3, we present examples from the Hateful Memes dataset, showcasing cases where the model made both correct and incorrect predictions.
In Figure 3(a), the Gold explanation describes the image as reinforcing a harmful racial stereotype by juxtaposing a joyful scene of Asian individuals eating with offensive text. The Predicted explanation correctly identifies the derogatory language and its racist implications, aligning with the gold annotation. The model’s BERT-F1 score of 0.873 shows the high confidence in associating textual and visual elements to detect hate speech effectively.
In Figure 3, the Gold explanation interprets the image as a humorous juxtaposition, using wordplay between nationality and species without targeting any group. However, the Predicted explanation classifies it as hateful. This missclassification suggests that the model struggled to distinguish linguistic humor from implicit hate speech, as reflected in its BERT-F1 score of 0.6259. This highlights the challenge of detecting context-dependent content, where intent and interpretation play a crucial role in classification.
D.2 Effect of Single vs. Multistage Training
In Figure 4 we notice that multi-stage model correctly predicts the label and generates a faithful explanation, identifying how the combination of the image and text reinforces gender stereotypes. In contrast, the single-stage model label the meme with incorrect label – not-propaganda. Figure 5 shows another example from the hateful memes dataset. The multi-stage model correctly labels the image as hateful and grounds its explanation in the offensive combination of the swastika and the mocking caption. Whereas the single-stage model misclassifies the image as not-hateful. This highlights the effectiveness of multistage optimization in aligning predictions with grounded reasoning.
(a) Correct prediction.
(b) Incorrect prediction.
Figure 3: Example of correct and incorrect label prediction with explanation.
Figure 4: Single- vs. Multi-Stage prediction for ArMeme.
Figure 5: Single- vs. Multi-Stage prediction for hateful memes.
Appendix E Data Release
The MemeXplain dataset222anonymous.com will be released under the CC BY-NC-SA 4.0 – Creative Commons Attribution 4.0 International License: https://creativecommons.org/licenses/by-nc-sa/4.0/.